The energy stored in a copper wire of cross sectional area 0.5mm² is 0.04J. calculate the force applied to it if its length is 2m. Take young modulus of copper as 1.2×10¹¹Nm^-2.

1
Expert's answer
2022-03-01T15:17:40-0500

"F=\sqrt{2kA}=\sqrt{\frac{2ESA}l}=49~N."